How to find the general solution of tanxtan4x=1? Please explain with steps

Grade:11

1 Answers

Arun
5 years ago
Dear Purva
 
tan 4x = 1 / tanx
 
tan 4x = cot x
 
tan 4x = tan (pi/2 – x)
 
4x = n * pi + (pi / 2 – x)
 
5x = n * pi + pi/2
 
x = 1/5 [n * pi + pi/2]